OpenStudy (anonymous):

Find the x-, y-, and z-intercepts for the equation. 9x + 3y - 6z = 18

OpenStudy (amistre64):

18/n where n is the coeff of the intercept your looking for gives you that intercept

OpenStudy (amistre64):

18/9,0,0 0, 18/3, 0 0,0 18/-6

OpenStudy (anonymous):

\[z = -3 + (3 x)/2 + y/2 \]\[9 x+3 y-6 z-18 = 0 \]\[3 (3 x + y - 2 z) = 18 \]

OpenStudy (anonymous):

There more comming

OpenStudy (anonymous):

\[Reduce[3 x + y = 6 + 2 z, {x, y, z}, \mathbb{Z}s] \]

OpenStudy (anonymous):

\[[3 x + y = 6 + 2 z, z] \]your answer Solve